OpenStudy (anonymous):

In a right triangle ABC, AB, the side opposite the 30 degrees angle measures 12 inches. Which is the length of the BC, the side opposite the 90 degrees angle? Answer 24 12

OpenStudy (anonymous):

other answers 2 times the square root of 24

OpenStudy (anonymous):

12/sin(30)

OpenStudy (anonymous):

so the answer is 12

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\sin30=12/x\] \[x=12/\sin30\] \[x=24\] sin30=1/2 12divided by 1/2=24

OpenStudy (anonymous):

24 because the short side must be half of the hypotenuse. no trig needed

OpenStudy (anonymous):

double the triangle and you will have an equilateral triangle with all sides 12 (equilateral because all degrees are 60) so it is obvious that the short side is half of the hypotenuse